Jan. 31 – Feb. 2

– Three stores provided ~44.25 hours of luminosity
– Kicker problems caused a loss of anti-proton transfers to the Recycler
– FESS found and later fixed an air conditioning problem in MI-31
– H Source had a problem with extraction
– Control fixed a communication problem with the A0 compressor room
– MI conducted safety system tests
– Linac worked on LRF3
– Power glitches due to the storm took down many RF systems. This caused the loss of the stash and MI interlocks to drop.
– Recovery operations took place due to snowstorm problems

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts